3. c) A wire of length 7 m and radius 2.37 mm is stretched by a load of 12370 gm and increased by 1.78% of the original length. Find the stress, strain, and Young's modulus of the material of the wire.

1
Expert's answer
2021-12-06T09:46:43-0500

Find the stress:


"\sigma=\frac{F}{A}=\frac {mg}{\pi r^2}=6.87\text{ MPa}."


The strain:


"\epsilon=1.78/100=0.0178."

Young's modulus of the material:


"E=\frac{\sigma}{\epsilon}=386\text{ MPa}."